A 1,000 kg car running at 50 km/h eastward a head-on collision with a 2,000 kg truck running at 30 km/h. If the two vehicles are locked together after collision, calculate the final velocity of the wreckage. What is the loss in kinetic energy of the system?


1
Expert's answer
2021-09-29T17:51:12-0400

We can find the final velocity of the wreckage from the law of conservation of energy:


"m_1v_{1i}-m_2v_{2i}=(m_1+m_2)v_f,""v_f=\dfrac{m_1v_{1i}-m_2v_{2i}}{m_1+m_2},""v_f=\dfrac{1000\ kg\cdot13.89\ \dfrac{m}{s}-2000\ kg\cdot8.33\ \dfrac{m}{s}}{1000\ kg+2000\ kg}=-0.92\ \dfrac{m}{s}."

The sign minus means that the final velocity of the wreckage directed to the west.

We can find the loss in kinetic energy of the system as follows:


"\Delta KE=KE_i-KE_f,""\Delta KE=\dfrac{1}{2}m_1v_{1i}^2+\dfrac{1}{2}m_2v_{2i}^2-\dfrac{1}{2}(m_1+m_2)v_{f}^2,"

"\Delta KE=\dfrac{1}{2}\cdot1000\ kg\cdot(13.89\ \dfrac{m}{s})^2+\dfrac{1}{2}\cdot2000\ kg\cdot(8.33\ \dfrac{m}{s})^2-\dfrac{1}{2}\cdot(1000\ kg+2000\ kg)\cdot(0.92\ \dfrac{m}{s})^2=165\ kJ."

Answer:

"v_f=0.92\ \dfrac{m}{s}," to the west.

"\Delta KE=165\ kJ."